# CCaaLF: Concurrency Control as a Learnable Function

### v3

@@ 2.3. State-of-the-Art Concurrency Control Algorithms
Table 1 compares state-of-the-art concurrency control algorithms within our design space. We omit timeout from the table because it is supported by all algorithms. It can be seen that some algorithms, such as the classic OCC and 2PL, support only one type of CC actions. Some recent algorithms, such as Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) and ASOCC (DBLP:conf/icde/LinCZ18/ASOCC, ), support multiple CC actions. It allows them to fine-tune conflict handling actions based on the workload and extract more concurrency from running transactions.
@@
The table reveals several gaps in the design space, i.e., there are CC actions not covered by existing algorithms. In particular, no algorithms consider priority-based conflict resolution in combination with detecting critical conflicts. Most algorithms, including the ones that support multiple CC actions, do not consider transaction priorities. CCaaLF introduces a novel learning-based approach for concurrency control, enabling support for all CC actions in the design space.
3. CCaaLF Overview
CCaaLF is a novel learned concurrency control algorithm that achieves high performance across different workloads and is efficient to optimize. Figure 3 shows the main components of CCaaLF: a collector that tracks real-time database states, a feature selector $\mathcal{{E}}$ that translates the collected states into a state vector $\mathbf{{s}}$, an in-database cache of an agent function $\mathcal{{F}}$ in the form of an action lookup table, and an optimizer that detects workload drifts and optimizes the agent function accordingly. We list the notations in Table 2.
CCaaLF operates across two workflows: the CC execution workflow and the agent function optimization workflow. The former is on the critical path of transaction execution and is invoked before every data access operation to determine the optimal concurrency control actions. This workflow takes the current operation $op$ and transaction $T$ as input, and outputs a vector of concurrency control actions $\mathbf{{a}}$ for the transaction manager to execute. Specifically, it consists of three steps. First, the transaction manager sends $(T,op)$ to the collector, which extracts database states based on $T$ and $op$ and combines them into a feature vector $\mathbf{{x}}$. Next, the feature selector $\mathcal{{E}}$ selects important features and generates a state $\mathbf{{s}}$, i.e., ${\mathbf{{s}}}=\mathcal{{E}}(\mathbf{{x}})$. Finally, the concurrency control actions $\mathbf{{a}}$ are retrieved from the lookup table, i.e., $\mathbf{{a}}=\mathcal{{F}}(\mathbf{{s}})$. This workflow is efficient enough such that it can be executed at per-operation granularity. We describe it in more detail in Section 4.
Table 2. Symbols and descriptions
$T$ the currently executing transaction
$op$ the currently executing transaction operation
${\mathbf{{x}}}$ the vector with all features
${\mathbf{{s}}}$ the database state
${\mathbf{{a}}}$ the CC actions for the current operation
${\mathcal{{F}}}$ the agent function
${\mathcal{{E}}}$ the feature selector
$\mathcal{{F}}$) the system performance achieved using $\mathcal{{F}}$
The optimization workflow is off the critical path of transaction execution. It maintains system performance under workload drifts, by monitoring the system performance and reoptimizing the action lookup table when necessary. In this workflow, data pairs of the current function $\mathcal{{F}}$ and the system performance $\text{Score}(\mathcal{{F}})$ are collected periodically. The data is analyzed to detect performance degradation caused by workload drifts. When workload drift is detected, CCaaLF starts optimizing the agent function for the current workload. This optimization process runs with a pre-defined time budget $t_{max}$. In the end, the newly optimized $\mathcal{F}_{next}$ is loaded into the database cache as a new action lookup table. We describe this workflow in more detail in Section 5.
4. CCaaLF Execution
In this section, we describe the CC execution workflow in CCaaLF, which lies on the critical path of transaction execution. We first present a model of concurrency control as a learnable function $\mathcal{{F}}$ that maps current database states to concurrency control actions. Next, we discuss how to efficiently capture the input of $\mathcal{{F}}$ and evaluate it quickly. Finally, we present the detailed algorithm steps.
4.1. Concurrency Control As Learnable Function
@@
Concurrency control algorithm can be modeled as a function $\mathcal{{F}}$ that maps the current database state $\mathbf{{s}}$ to a combination of conflict detection and conflict resolution actions $\mathbf{{a}}$:
@@
Before executing an operation, CCaaLF captures current database states $\mathbf{{s}}$, computes $\mathcal{{F}}$($\mathbf{{s}}$), and executes the resulting actions $\mathbf{{a}}$. After all operations are executed, CCaaLF starts a validation process before committing the transaction. This validation process detects and resolves previously ignored conflicts, and aborts the transaction when the conflicts cannot be resolved, for example, when the transaction reads a data object from an aborted transaction.
@@
The function $\mathcal{{F}}$ covers many existing CC algorithms, including Polyjuice, ASOCC, 2PL, LSDF, OCC, and IC3. We demonstrate it with two examples of 2PL (wait die) and ASOCC. Let $T_{req}$ and $T_{owner}$ be the timestamps of the current transaction and of the oldest lock owner. Let $hotness$ denote the hotness level of the data object being accessed. 2PL and ASOCC can be modeled as follows:
$\mathcal{F}_{\text{2PL}}(T_{req},T_{owner})=\left\{\begin{aligned} (\text{{detect all}},0,\text{FIFO pri.})&&T_{req}>T_{owner}\\ (\text{{detect all}},\infty,\text{FIFO pri.})&&\text{otherwise}\\ \end{aligned}\right.$
$\mathcal{F}_{\text{ASOCC}}(hotness)=\left\{\begin{aligned} (\text{{no detection}},0,\text{No pri.})&&hotness=\text{cold}\\ (\text{{detect all}},\infty,\text{No pri.})&&hotness=\text{hot}\\ (\text{{detect critical}},0,\text{No pri.})&&\text{otherwise}\\ \end{aligned}\right.$
@@
Each action consists of a conflict detection technique, a timeout value, and a transaction priority (the no priority option means all transactions have the same priority). In these examples, $T_{req}$, $T_{owner}$, and $hotness$ constitute the input states $\mathbf{s}$ of 2PL and ASOCC, respectively. The function $\mathcal{F}_{\text{2PL}}$ and $\mathcal{F}_{\text{ASOCC}}$ output actions which are then executed by the CC algorithm. In particular, 2PL aborts the transaction when $T_{req}$ is higher than $T_{owner}$, otherwise, it blocks the transaction and wakes it up later in a FIFO order. ASOCC ignores conflicts for cold data. It detects all conflicts and blocks the transaction for hot data. In other cases, it aborts the transaction when it finds critical conflicts.
4.2. State Capturing
Figure 4. CCaaLF’s state capturing process.
@@ 4.2. State Capturing
The input $\mathbf{{s}}$ of $\mathcal{{F}}$ is important as it determines the quality of CC actions, and it contributes to the CC algorithm’s overhead. There are two challenges in capturing $\mathbf{{s}}$. First, it must be done quickly since it is on the critical path of transaction execution. Modern in-memory databases can commit a transaction in microseconds, therefore $\mathbf{{s}}$ should be captured in order of nanoseconds. Second, the input must contain relevant database states so that the optimization of CC actions can extract a high degree of concurrency. Furthermore, the input space must be small to minimize storage overhead and to speed up the optimization of $\mathcal{{F}}$.
Figure 4 shows the state-capturing process in CCaaLF. It employs two optimizations to address the challenges above. The first optimization is to use a fixed set of lightweight features that can be efficiently captured across transactions in a lock-free manner. CCaaLF uses nine features, grouped into transaction features and graph features, as shown in Figure 4. These features have been shown to be effective at capturing transaction and dependency graph characteristics (DBLP:journals/pvldb/TianHMS18/LDSF, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/sigmod/HuangMSW17/VATS, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ). More importantly, they are common features readily available in transactional databases, and therefore are not specific to any concurrency control algorithm. The second optimization leverages a feature selector $\mathcal{{E}}$ that removes non-influential features. The result is a small set of states that is used as the input for $\mathcal{{F}}$. The selected features achieve a balance between the quality of the optimized agent function and the optimization time. CCaaLF provides a default feature selector that achieves good performance. The user can improve this default selector, at some cost, by running the optimization algorithm described in Section 5.5.
Figure 5. CCaaLF’s action lookup table.
4.3. Function Evaluation
@@
The challenge in designing $\mathcal{{F}}$ is to ensure both efficiency and effectiveness. Specifically, the function must be quick to compute, since it is on the transaction’s critical path. Furthermore, the output actions must be able to extract high degrees of concurrency across different workloads. For efficiency, CCaaLF implements the function as a learnable action lookup table, such that the function computation reduces to a single table lookup. Although implementing $\mathcal{{F}}$ as complex machine learning models can approximate the state-action mapping more accurately, the inference cost is too high for concurrency control. In contrast, by learning the agent function in the form of a lookup table, the actions can be computed in only hundreds of CPU cycles.
@@
To achieve high performance across diverse workloads, $\mathcal{{F}}$ learns combinations of design choices from existing CC algorithms optimized for different workloads. An output of $\mathcal{{F}}$ contains a conflict detection mechanism, which can be either no detection, detect critical, or detect all. It also contains a timeout value and a transaction priority for conflict resolution. Figure 5 shows an action lookup table where the first column is the input state and the other columns make up the CC actions. We will discuss the columns in more detail in Section 4.5. CCaaLF implements several optimizations for the agent function, including pruning suboptimal functions from the search space, learning transaction backoffs, and transforming some actions into more learnable forms.
Assign high priority after detect all. Unlike 2PL, in CCaaLF, it is possible for an operation that passes the detect all check to be ordered after another operation with a higher priority, leading to its eventual abort and retry even if both of them are using detect all action. Thus, we heuristically assign the highest priority to transactions that have passed the detect all check, avoiding such unnecessary aborts.
Validate before expose. We notice that exposing dirty data must be done carefully, especially when combined with no detect actions, as aborting a transaction that has exposed its data can result in a long chain of cascading aborts. Thus, we follow existing work (DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/eurosys/MuAS19/DRP, ; DBLP:conf/sosp/XieSLAK015/Callas, ) to perform an early validation before exposing dirty data.
Learn transaction backoff. We follow existing studies (DBLP:conf/sigmod/ShengTZP19/abort_prediction, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ) to learn a backoff for transaction abort and retry. This number is learned together with timeout, for each type of transaction.
4.4. Function Update
The agent function is updated by optimizing a new function to replace the existing one when workload drifts are detected. CCaaLF employs a simple, yet effective threshold-based method for workload drift detection. Specifically, it monitors system performance during database operations and triggers the optimization process upon detecting significant performance changes. The threshold for detection is 10% (in throughput). This value is determined empirically to ensure robust detection and to avoid over-triggering. We will show that our detection method is effective in Section 6.2. We emphasize that CCaaLF only focuses on the design of the concurrency control algorithm. Thus, the problem of workload drift detection is orthogonal. In other words, a detailed analysis of how our detection method compares to other algorithms, such as those discussed in  (DBLP:journals/pacmmod/WuI24/2sample, ), is beyond the scope of this paper.
CCaaLF enables agent function updates without blocking the ongoing or incoming transactions. In particular, each transaction is assigned a specific agent function, namely the latest one, at the start of its execution. During an agent function update, CCaaLF maintains both the latest function and the previously loaded functions still in use by ongoing transactions. While a new function is being loaded, the old agent function continues to handle transactions. Once the loading is complete, the system performs an atomic pointer update to switch to the new agent function. This design avoids blocking ongoing or incoming transactions. It also ensures agent function consistency, that is, it prevents the switch of functions in the middle of a transaction’s execution.
4.5. Detailed Algorithm
We describe how CCaaLF executes an operation based on the learned actions. Given an operation $op$ of transaction $T$, the algorithm first determines the CC actions using the learned function implemented by the GetCC function defined below.
(2) $GetCC(T,op)$ $\mapsto\mathbf{{a}}=\langle\mathbf{{a}}_{d},\mathbf{{a}}_{t},\mathbf{{a}}_{p},(w_{1},\dots,w_{n},\textit{expose})\rangle$
@@ 4.5. Detailed Algorithm
$\mathbf{{a}}_{d}$, $\mathbf{{a}}_{t}$, and $\mathbf{{a}}_{p}$ are the conflict detection method, timeout and transaction priority, respectively. $(w_{1},\dots,w_{n})$ represents the rule for identifying critical conflicts (as explained in Section 2.1). Specifically, let $n$ be the number of transaction types, $w_{i}$ means that $T$ considers the first $w_{i}$ operations of a dependent transaction of type $i$ as critical conflicts. The boolean parameter $expose$ determines whether $T$ makes its operations visible to other transactions after executing $op$.
1 Function Execute($T,op$):
2       a $\leftarrow$ GetCC($T,op$)

### v4

@@ 2.3. State-of-the-Art Concurrency Control Algorithms
Table 1 compares state-of-the-art concurrency control algorithms within the design space. We omit timeout from the table because it is supported by all algorithms. It can be seen that some algorithms, such as OCC and 2PL, support only one type of CC actions. Other algorithms, such as Polyjuice (Wang et al., 2021) and ASOCC (Lin et al., 2018), support multiple CC actions, with which the algorithm can fine-tune conflict handling actions based on the workload to extract more concurrency.
@@
The table reveals several gaps in the design space, i.e., CC actions that are not covered by existing algorithms. For example, no existing algorithms consider priority-based conflict resolution in combination with detecting critical conflicts. Most algorithms, including the ones that support multiple CC actions, do not consider tuned transaction priorities.
3. NeurCC Overview
In this section, we present an overview of NeurCC, a novel learning-based concurrency control algorithm that achieves high performance across diverse workloads. It supports the complete set of CC actions shown in Table 1, by capturing existing concurrency control algorithm designs via a learnable function. In the following, we describe the learning function, and the workflow that enables its efficient execution and optimization. We summarize the notations in Table 2.
3.1. Concurrency Control As a Learnable Function
Table 2. Summary of notations
$T$ current transaction
opop current operation of the current transaction
$\mathbf{{x}}$} vector with all features
$\mathbf{{s}}$} database state
$\mathbf{{a}}$} CC actions for the current operation
$\mathcal{{F}}$} function
$\mathcal{{E}}$} feature selector
Score($\mathcal{{F}}$) system performance using $\mathcal{{F}}$
@@ 3.1. Concurrency Control As a Learnable Function
We model concurrency control algorithm as a function $\mathcal{{F}}$ that maps the current database state $\mathbf{{s}}$ to a combination of conflict detection and conflict resolution actions $\mathbf{{a}}$.
@@
The state $\mathbf{{s}}$ comprises features used in various CC algorithms, such as transaction access ID, transaction type, and data hotness. The actions $\mathbf{{a}}$ include conflict detection actions, timeout values, and pipeline-wait actions. Before executing an operation, the algorithm collects the current database state $\mathbf{{s}}$, computes $\mathcal{{F}}$($\mathbf{{s}}$), and then executes the resulting actions $\mathbf{{a}}$.
@@
$\mathcal{{F}}$ captures many existing CC algorithms, including Polyjuice, ASOCC, 2PL, bLDSF, OCC, and IC3. We demonstrate this with two examples of 2PL (wait-die) and ASOCC. Let $t_{\textit{req}}$ and $t_{\textit{owner}}$ be the timestamps of the current transaction and of the oldest lock owner. Let hotnesshotness be the hotness level of the data object being accessed. 2PL (wait-die) and ASOCC can be modeled as follows:
[…]
@@
Each action consists of a conflict detection action, a timeout value, and a transaction priority ($\varnothing$ represents no priority). In these examples, $t_{req}$, $t_{owner}$, and hotnesshotness are the state $\mathbf{s}$. The output of $\mathcal{F}_{\text{2PL}}$ and $\mathcal{F}_{\text{ASOCC}}$ are then executed by the algorithm. In particular, 2PL (wait-die) aborts the transaction when $t_{\textit{req}}$ is higher than $t_{\textit{owner}}$; otherwise, it blocks the transaction and wakes it up later in the timestamp order. ASOCC ignores conflicts for cold data. It detects all conflicts and blocks the transaction for hot data. In other cases, it aborts the transaction when there are critical conflicts.
We define the problem of finding $\mathcal{{F}}$ that extracts the most concurrency across diverse workloads as an optimization problem. Let $\mathbf{S}$ be the set of possible input states $\mathbf{{s}}$, Score($\mathcal{{F}}$) be the system performance [^1] We use throughput as the performance metric, similar to previous works (Wang et al., 2021; Tang et al., 2017). using $\mathcal{{F}}$, and $\mathbf{A}$ be the set of possible output actions $\mathbf{{a}}$. The optimization objective is defined as:
(2) maximize $\text{Score}(\mathcal{{F}})$
NeurCC starts the optimization process when it detects a workload drift. Specifically, there is a background worker, named optimizeroptimizer, that continuously monitors the system’s current throughput, i.e. the current number of committed transactions per second (TPS). The optimizeroptimizer tracks the throughput regardless of whether the system is saturated. Optimization is triggered when the relative change in throughput exceeds a predefined threshold. In the current design, the threshold value is determined empirically, e.g., $10\%$. The value is chosen to achieve reliable detection while avoiding unnecessary optimization triggers. We discuss this threshold in Section 7.2, and provide more details in Section 4.3. We note that more complex workload drift detection techniques (e.g., two-sample-test (Wu and Ives, 2024)) could also be used. Since our main focus is on the learning of concurrency control, we leave the design and integration of advanced drift detectors to future work.
3.2. NeurCC Workflow
Figure 3. NeurCC’s concurrency control (CC) execution workflow and function optimization workflow.
[…]
Figure 3 shows the main components of NeurCC: a collector that tracks real-time database states, a feature selector $\mathcal{{E}}$ that maps the collected states into a state vector $\mathbf{{s}}$, an in-database cache that maintains the function $\mathcal{{F}}$ in the form of an action lookup table, and an optimizer that detects workload drifts and optimizes the function upon detection.
NeurCC operates across two workflows: the CC execution workflow and the function optimization workflow. The former is on the critical path of transaction execution. It is invoked before every data access operation to select the optimal actions. This workflow takes the current operation $op$ and transaction $T$ as input, and outputs a vector of actions $\mathbf{{a}}$ for the transaction manager to execute. The transaction manager first sends $(T,op)$ to the collector, which collects database states based on $T$ and $op$, then combines them into a feature vector $\mathbf{{x}}$. Next, the feature selector $\mathcal{{E}}$ selects important features from $\mathbf{{x}}$ and generates a state $\mathbf{{s}}$. Finally, the concurrency control actions $\mathbf{{a}}$ are returned from the lookup table, i.e., $\mathbf{{a}}=\mathcal{{F}}(\mathbf{{s}})$. This workflow is efficient such that it can be executed at per-operation granularity. We describe it in more detail in Section 4.
The second workflow is off the critical path of transaction execution. It maintains high performance under workload drifts by monitoring the system performance and optimizing the action lookup table when necessary. In this workflow, the function $\mathcal{{F}}$ and the corresponding system performance $\text{Score}(\mathcal{{F}})$ are collected periodically. The data is analyzed to detect workload drifts based on performance changes. When a drift is detected, NeurCC starts optimizing the function for the current workload. The optimization process runs with a pre-defined time budget $t_{\textit{max}}$, at the end of which the new function $\mathcal{F}_{\textit{next}}$ is loaded into the database cache as a new action lookup table. We describe this workflow in more detail in Section 5.
3.3. Discussion
NeurCC evaluates the agent function at per-operation granularity. This is because important CC mechanisms, such as dirty-read visibility and pipeline waits, work at the level of individual operations (as opposed to at the level of individual transaction). We note that coarse-grained decisions can be beneficial, for example, in enabling scheduling operations before execution (as in deterministic CC). Integrating such coarse-grained decisions with our fine-grained agent function is left as future work.
The current design of NeurCC targets CC paradigms based on per-operation conflict handling, which keeps the state and action space of the agent function small and efficient to optimize. We note that NeurCC captures some transaction-chopping behaviors through its pipeline-wait actions (i.e., $w_{i}$ and exposeexpose), and it can reproduce the behavior of algorithms like IC3 (Wang et al., 2016) that uses transaction chopping. Other paradigms, such as deterministic CC, sagas, or timestamp-ordering, entail more complex state and action representations (e.g., actions that schedule transactions before execution). Supporting them would significantly expand the agent function’s state and action space, thus requiring new learning algorithms, which we leave for future work.
4. NeurCC Execution
In this section, we describe NeurCC’s execution workflow which lies on the critical path of transaction execution. We first discuss what database states NeurCC collects, and how it selects important states as input to the function $\mathcal{{F}}$. We then present the detailed concurrency control algorithm that performs conflict handling based on the output of $\mathcal{{F}}$.
4.1. State Collection and Selection
@@
Before executing a transaction operation, NeurCC first computes a vector $\mathbf{s}$ that captures the current, relevant database states (Figure 4). The vector $\mathbf{{s}}$ is important as it determines the quality of CC actions generated by $\mathcal{{F}}$. There are two challenges in computing $\mathbf{{s}}$. First, it must be efficient since it is on the critical path of transaction execution. Modern in-memory databases can commit a transaction in microseconds, therefore $\mathbf{{s}}$ should be computed in order of nanoseconds. Second, $\mathbf{{s}}$ must contain relevant database states so that CC actions returned by $\mathcal{{F}}$ can extract a high degree of concurrency, while remaining small to minimize storage overhead and to speed up the optimization of $\mathcal{{F}}$.
Table 3. Features collected by NeurCC. $T$ denotes the current transaction.
Feature Definition
# of executed SQLs Number of SQLs already executed by $T$.
Read dirty data If $T$ has read uncommitted dirty data from others.
Transaction type Identifier of the transaction template to which $T$ belongs (e.g., NewOrder, Payment).
Access ID Identifier of the currently executing SQL.
Operation type The current operation type (read/write).
Data hotness Contention level of the accessed record (cold, warm, hot), derived from data access frequency and data correlation (Lin et al., 2018).
# of dependent txns. Number of transactions $T$ depends on.
# of transactions Number of running transactions.
Out-degree Number of transactions that depend on $T$.
NeurCC employs two techniques to address the challenges above. The first is to use a fixed set of lightweight states (or features) that can be efficiently captured across transactions, in a lock-free manner. NeurCC collects nine features, which are listed in Table 3 and can be grouped into transaction features and graph features. These features have been shown to be effective at capturing transaction and dependency graph characteristics (Tian et al., 2018; Lin et al., 2018; Wang et al., 2021, 2016; Huang et al., 2017a; Su et al., 2017). More importantly, they are common features readily available in transactional databases, that is, they are not specific to any concurrency control algorithm.
The second technique is to use a feature selector $\mathcal{{E}}$ that selects influential features to achieve a balance between the quality of the optimized function and the optimization time. NeurCC provides a default $\mathcal{{E}}$ that achieves good performance, which the user can improve at some cost, by running the optimization algorithm described in Section 5.4.
Figure 4. NeurCC’s state collection and selection process.
[…]
Before a transaction accesses a data record, NeurCC collects the current database state as a vector $\mathbf{{s}}$. Consider ASOCC as a running example. This vector includes data hotness feature, thus access to records with different contention levels is mapped to different states, which enables the selection of appropriate conflict-handling actions as in ASOCC.
4.2. Action Function
@@
After computing $\mathbf{{s}}$, NeurCC executes $\mathcal{{F}}$($\mathbf{{s}}$) to get the concurrency control actions $\mathbf{{a}}$ for the current transaction operation. The main challenge in designing $\mathcal{{F}}$ is to ensure both efficiency and effectiveness. Specifically, the function must be quick to evaluate, since it is on the transaction’s critical path. Furthermore, the output actions must be able to extract high degrees of concurrency across different workloads. For efficiency, NeurCC implements $\mathcal{{F}}$ as a learnable action lookup table, reducing the function evaluation to a single table lookup. Although implementing $\mathcal{{F}}$ as a machine learning model can approximate the state-action mapping more accurately, the inference cost is too high for concurrency control. In contrast, the lookup table can return the actions in hundreds of CPU cycles.
Figure 5. NeurCC function implemented as a lookup table.
[…]
@@
To achieve high performance across diverse workloads, $\mathcal{{F}}$ learns combinations of design choices from existing CC algorithms optimized for different workloads. An output of $\mathcal{{F}}$ contains a conflict detection mechanism, which can be either no detection, detect critical, or detect all. It also contains a timeout value and a transaction priority for conflict resolution. Figure 5 shows an action lookup table, where the first column is the input state and the remaining columns make up the CC actions. Let $n$ denote the number of transaction types. The table has the following schema:
(4) $\langle\mathbf{{a}}_{d},\mathbf{{a}}_{t},\mathbf{{a}}_{p},(w_{1},\dots,w_{n},\textit{expose})\rangle$
@@
$\mathbf{{a}}_{d}$, $\mathbf{{a}}_{t}$, and $\mathbf{{a}}_{p}$ are the conflict detection method, timeout, and transaction priority, respectively. $(w_{1},\dots,w_{n})$ represents the rule for identifying critical conflicts (as described in Section 2.1). Specifically, $w_{i}$ means that $T$ considers the first $w_{i}$ accesses of a dependent transaction with type $i$ as critical conflicts, thereby preventing the formation of dependency cycles among concurrent transactions that might read dirty data from each other. exposeexpose determines if $T$ makes its accesses visible to other transactions after executing $op$.
In the ASOCC example, NeurCC selects no detection for cold data, which corresponds to optimistic execution. It selects detect all for hot data, forcing conflict detection for operations accessing highly contended records. It selects detect critical for warm data, enabling early validation to proactively abort transactions that are likely to fail.
NeurCC implements several optimizations when maintaining this action lookup table, including pruning suboptimal functions from the search space and learning transaction backoff strategies.
Assign high priority after detect all. Unlike 2PL, in NeurCC, it is possible for an operation that passes the detect all check to be ordered after another operation with a higher priority, leading to its eventual abort and retry even if both of them are using detect all action. Thus, we heuristically assign the highest priority to transactions that have passed the detect all check, avoiding such unnecessary aborts.
Validate before expose. We notice that exposing dirty data must be done carefully, especially when combined with no detect actions, as aborting a transaction that has exposed its data can result in a long chain of cascading aborts. Thus, we follow existing work (Wang et al., 2021, 2016; Mu et al., 2019; Xie et al., 2015) to perform an early validation before exposing dirty data.
Learn transaction backoff. We follow existing studies (Sheng et al., 2019; Wang et al., 2021) to learn a backoff for transaction abort and retry. This number is learned together with timeout, for each type of transaction.
4.3. Function Update
The function is updated by optimizing a new function to replace the existing one when workload drifts are detected. NeurCC employs a simple, yet effective threshold-based method for workload drift detection. Specifically, it monitors system performance and triggers the optimization process when detecting significant performance changes. The threshold for detection is 10% (in throughput), which is determined empirically to ensure robust detection and to avoid over-triggering. Section 7.2 shows that this detection method is effective. We emphasize that NeurCC only focuses on the design of the concurrency control algorithm, therefore the problem of workload drift detection is orthogonal. A detailed analysis of how our detection method compares to other algorithms, such as those discussed in  (Wu and Ives, 2024), is beyond the scope of the paper.
NeurCC performs function updates without blocking the ongoing or incoming transactions. In particular, each transaction is assigned a specific function, namely the latest one, at the start of its execution. During a function update, NeurCC maintains both the latest function and the previously loaded functions still in use by ongoing transactions. While a new function is being loaded, the old function continues to handle transactions. Once the loading is complete, the system performs an atomic pointer update to switch the latest function to the newly loaded one. This design avoids blocking ongoing or incoming transactions, and prevents switching functions in the middle of a transaction’s execution.
4.4. Transaction Execution
After determining the CC actions, NeurCC proceeds to execute the transaction operations. Similar to OCC, NeurCC does not guarantee all conflicts are detected and resolved during execution. Instead, it performs a validation before committing the transaction, once all operations are executed. We elaborate on these steps below.
1Function Execute($T,op$):
2  a $\leftarrow$ GetCC($T,op$) // compute $\mathbf{{s}}$ and $\mathcal{{F}}$($\mathbf{{s}}$)